# Combinatorial Selection with Costly Information

### v1

@@ 6 The Weighing Scale Problem (Minimization)
We will now introduce the Weighing Scale (henceforth, WS) problem. A decision maker is presented with $n$ alternatives $(X_{i},c_{i})$ and a combinatorial constraint $\mathcal{F}\subseteq 2^{[n]}$; $X_{i}\geq 0$ is the random value of the alternative realized independently by a known distribution and $c_{i}\geq 0$ is a weighing cost. The only way the decision maker can determine any further information about each value $X_{i}$ beyond its distribution, is to use a weighing scale to compare it against some fixed threshold $t$ of their choosing at the additional cost of $c_{i}$ and learn whether $X_{i}\leq t$ or not. The process terminates with the decision maker selecting a feasible set of alternatives $S\subseteq\mathcal{F}$, paying their total value.
Note that each alternative $(X_{i},c_{i})$ corresponds to a Costly Information MDP $\mathcal{M}_{i}$ that captures the information acquisition process described above. In particular, the available actions at the starting state of each MDP $\mathcal{M}_{i}$ are as follows:
@@
    Pick a threshold $t\in\mathrm{support}(X_{i})$ and weigh the alternative against it at a cost of $c_{i}$. Upon taking one of these actions, the MDP advances to one of two random subprocesses $\mathcal{M}_{i}^{\leq t}$ and $\mathcal{M}_{i}^{>t}$, defined over the random variables $(X_{i}|X_{i}\leq t)$ and $(X_{i}|X_{i}>t)$ respectively, based on the random outcome of the weighing.
@@
From this equivalence, an instance of WS corresponds to an instance of min-CICS over the corresponding MDPs. [^10] Technically, our framework does not capture infinite horizon MDPs. However, observe that whenever the decision maker has identified that the value of the alternative lies in some interval of length $\leq c$, performing any extra weighings is suboptimal. Thus, the corresponding MDPs for the WS problem are finite horizon without loss. Our main result for this section is the following.
@@
For any instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ of matroid-WS, there exists an efficient algorithm that achieves an $O(\max_{i}\kappa_{i})$-approximation to the optimal adaptive policy, with the parameter $\kappa_{i}$ for each alternative $i\in[n]$ defined as
@@
We prove Theorem 6.1 by showing that for each alternative $i\in[n]$, the MDP $\mathcal{M}_{i}$ admits an $O(\kappa_{i})$-pointwise approximation. We note that without any assumptions on the distributions of the values, the parameters $\kappa_{i}$ can be unbounded; in particular, for heavy tail distributions, $\mu_{i}/M_{i}$ can be arbitrarily large. We show in Theorem 6.2 (proven in Appendix D) that this dependence is necessary for any pointwise approximation guarantee. Better bounds may be possible via the weaker notion of local approximation or a global argument.
@@
For any $\alpha\geq 1$, there exists a WS alternative $\mathcal{M}$ that does not admit an $\alpha$-pointwise approximation.
Proof of Theorem 6.1.
We begin by introducing a committing policy for WS described below, which we call the One-Sided Halving algorithm. The policy begins by weighing the alternative against some threshold $t_{2}$; if the alternative is larger, then the policy commits to no more weighings and if it is smaller, it halves its threshold from $t_{2}$ to $t_{2}/2$ and repeats, until the threshold reaches some fixed lower bound $t_{1}$.
@@
Input : An MDP $\mathcal{M}$ corresponding to an alternative $(X,c)$ and two thresholds $0<t_{1}<t_{2}$.
@@
Let $h$ be the solution to equation $c=\operatorname{E}\left[(X-h)^{+}\right]$. The following lemma, combined with our reduction from pointwise to local approximation (2) and our local approximation composition algorithm (Corollary 4.2), directly implies Theorem 6.1.
@@
The rest of this section is dedicated towards proving Lemma 6.3. We fix a WS alternative $\mathcal{M}$ corresponding to random value $X$ of support $\mathcal{X}:=\mathrm{support}(X)$ and weighing cost $c$. We use $\mu,M,g,h$ and $\kappa$ to denote the parameters of the alternative, as previously defined. In order to develop our pointwise approximation guarantees, we first need to obtain expressions for the surrogate costs.
Surrogate cost of $\mathcal{M}$.
Fact 3.
$\rho^{*}(x):=\min(h,\max(g,x)).$
Committing policies.
A committing policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ for WS will declare in advance a decision tree over pre-specified thresholds against which it will weigh $X$, resulting in a Markov chain $\mathcal{M}^{\pi}$. Importantly, $\pi$ will end up partitioning the support $\mathcal{X}$ of distribution $\mathcal{D}$ into a set of intervals, corresponding to its terminal states. Furthermore, the probability of running the Markov chain $\mathcal{M}^{\pi}$ and ending up in a terminal state $t$ corresponding to some interval $I$ will be precisely $\operatorname{Pr}\left[X\in I\right]$. A pictorial representation of such Markov chains is given in Figure 5. We use $\mathcal{I}^{\pi}$ to denote this set of intervals, and $t(I)$ to denote the terminal state of $\mathcal{M}^{\pi}$ corresponding to interval $I\in\mathcal{I}^{\pi}$. Furthermore, the probability of running the Markov chain $\mathcal{M}^{\pi}$ and ending up in a terminal state $t(I)$ will be precisely $\operatorname{Pr}\left[X\in I\right]$. This allows us to obtain the following characterization of surrogate costs for committing policies.
Fact 4.
$\rho^{\pi}(x):=W^{*}_{\mathcal{M}^{\pi}}(t(I))$
for the unique interval $I\in\mathcal{I}^{\pi}$ that contains $x$.
@@
$p(s)\cdot c(s)=\sum_{t\in T(S)}p(t)\cdot b_{st}$
$k:=\log\frac{t_{2}}{t_{f}}\leq\log\frac{M}{g}\leq\log\frac{2\mu}{g}$
$W^{*}_{\mathcal{M}^{\pi}}(t(I))\leq\mu(I)+k\cdot g$
To summarize, we have shown that for all $x\in\mathcal{X}$:
$\rho^{\pi}(x)\leq u(x):=kg+2\cdot\min(2\mu,\max(x,g)).$
Notice that the upper bound $u(x)$ is non-decreasing. Now, recall that $\rho^{*}(x)=\min(h,\max(g,x))$ is also a non-decreasing mapping. Thus, policy $\pi$ will $\alpha$-pointwise approximate $\mathcal{M}$ for
$a=\max_{x\in X}\frac{u(x)}{\rho^{*}(x)}$
7 Pandora’s Box with Optional Inspection and Semi-Local Approximation
Finally, we consider Pandora’s Box with Optional Inspection (henceforth PBOI), a well-studied variant of Pandora’s Box in which a box may be selected without first being opened. For brevity, we call selecting a box without first opening it “grabbing” the box. In the cost minimization (resp. reward maximization) setting, if the decision maker grabs a box, they pay (earn) the cost (value) inside of the box immediately, without ever having observed it. The availability of this extra action means that each box is now an MDP instead of a Markov chain, leading to a significantly more challenging optimization problem.
Formally, we represent a box as a pair $\mathcal{B}=(\mathcal{D},c)$, where $\mathcal{D}$ is a distribution over nonnegative values, and $c>0$ is the opening cost. Figure 6 illustrates the actions available to the algorithm.
Figure 6: A Pandora’s box $\mathcal{B}=(\mathcal{D},c)$ under optional inspection is initially closed. In order to learn the value realization $X\sim\mathcal{D}$, the decision maker can open it (at a cost of $c$), after which they can select it to accept its value $X$. Alternatively, the decision maker can grab the box, meaning select it without opening it, receiving the box’s value $X$ without paying any cost. Because $X$ is unknown whenever the box is being grabbed, for the purposes of expected value, receiving $X$ is equivalent to receiving $\mathbf{E}[X]=m$.
Previous work and our results.
As we mentioned earlier, the maximization version of PBOI admits a PTAS in the single-item selection setting. However, prior to this work, there was no known policy that achieves better than a $1/2$-approximation for matroid selection PBOI. Additionally, this $1/2$-approximation is somewhat uninteresting. As Beyhaghi and Kleinberg (2019) observe, the optimal policy that does not grab any box will attain at least as much utility as the utility earned by the optimal algorithm from boxes that it opens, and the optimal policy that only grabs boxes will attain at least as much utility as the utility earned by the optimal algorithm from boxes it grabbed. Thus, the sum of these policy’s utilities is an upper bound on the optimal algorithm and so uniformly randomizing between these policies will immediately give us a 1/2-approximation. [^11] Beyhaghi and Kleinberg (2019) make this observation in the single-item setting, but combining it with results of Singla (2017) generalizes it to the matroid setting.
One might wonder whether one can use local approximation to break the $1/2$-approximation barrier in the matroid selection setting. Alas, in the maximization setting, we cannot: there exist boxes for which we cannot do better than a local $(1/2+\varepsilon)$-approximation for any $\varepsilon>0$ (Theorem 7.2). Our contribution is a novel approach to break the $1/2$-approximation barrier in the matroid selection setting, proving the following result.
@@ 7 Pandora’s Box with Optional Inspection and Semi-Local Approximation
There exists an efficient randomized committing policy, namely Algorithm 3 with appropriately chosen probabilities, that achieves a $0.582$-approximation to the optimal adaptive policy for any instance of max-matroid-PBOI.
The key idea behind Theorem 7.1 is a refinement of local approximation, which we call semilocal approximation (Definition 11). Just as local approximations compose, we show in Theorem 7.3 that semilocal approximations similarly compose. We then show in Theorem 7.7 that, unlike ordinary local approximation, all boxes admit a semilocal approximation that leads to a $0.582$-global approximation. Theorem 7.1 then follows immediately from Theorems 7.3 and 7.7.
@@
There are two prices we pay in refining local approximation to semilocal approximation. First, the definition as we state it currently is specific to PBOI, though it could perhaps be generalized to other MDP families in the future. Second, semilocal composition only holds for matroid selection problems, whereas past works showed results compatible with any frugal (roughly, “greedy”) algorithm (Singla, 2017; Gupta et al., 2019; Scully and Doval, 2024). We believe this second obstacle may be fundamental.
7.1 Notation and preliminaries
In this section, we at times consider Pandora’s box problems under both optional inspection, as illustrated in Figure 6, and mandatory inspection, which is the traditional setting (Weitzman, 1979) that lacks the grab action. We use the same $\mathcal{B}=(\mathcal{D},c)$ notation for both, with context making it clear whether we are considering the box to be optional or mandatory inspection. We will apply our amortization framework for max-CICS, which is similar to the framework established in Section 3 and Section 4, and is detailed in Appendix E.
For a box $\mathcal{B}=(\mathcal{D},c)$ with optional inspection, the optimality curve of the local game $(\mathcal{B},y)$ is given by $f_{\mathcal{B}}(y):=\max\{\mathbf{E}_{X\sim\mathcal{D}}[\max\{X,y\}]-c,\mathbf{E}[X],y\}$, with the quantities in the max representing the options of open, grab, or choosing the outside option, respectively. We can also write the optimality curve of the open action as $f^{\mathrm{open}}_{\mathcal{B}}(y):=\max\{\mathbf{E}_{X\sim\mathcal{D}}[\max\{X,y\}]-c,y\}$. These allow us to deduce the water draining surrogate values for the box. We define the following quantities of interest.
  • The mean value is $m=\mathbf{E}_{X\sim\mathcal{D}}[X]$.
  • The Gittins index is $g$, the solution to $c=\mathbf{E}_{X\sim\mathcal{D}}[(X-g)^{+}]$. Note that $g$ is the water draining index of the policy that commits to opening the box, in other words, the largest outside option for which opening the box is better than accepting the outside option.
  • The backup index is $h=\max\{m,h^{\prime}\}$, where $h^{\prime}$ is the solution to $c=\mathbf{E}_{X\sim\mathcal{D}}[(h^{\prime}-X)^{+}]$. Note that $h^{\prime}$ is the largest outside option for which grabbing the box is better than opening it, and $h$ is the largest outside option for which grabbing is better than both of the alternatives.
In terms of these quantities, we can rewrite the optimality curve for $(\mathcal{B},y)$ as:
$f_{\mathcal{B}}(y)=\begin{cases}m&\text{if $y<h$}\\ \operatorname{E}\left[\max(y,\min(g,X))\right]&\text{if $y\in[h,g]$}\\ y&\text{if $y>g$}\end{cases}$
Recall that the optimal surrogate value of the box, as well as the surrogate value for the opening action, are defined so that $f_{\mathcal{B}}(y)=\mathbf{E}[\max\{W^{*},y\}]$ and $f^{\mathrm{open}}_{\mathcal{B}}(y)=\mathbf{E}[\max\{W^{\mathrm{open}},y\}]$. We therefore obtain:
  • The opening surrogate value is $W^{\mathrm{open}}=\min\{X,g\}$, where $X\sim\mathcal{D}$ is the box’s hidden value.
  • The optimal surrogate value is $W^{*}=\max\{\min\{X,g\},h\}=\max\{W^{\mathrm{open}},h\}$, where $X\sim\mathcal{D}$ is the box’s hidden value.
When the box in question is not clear from context, we clarify using subscripts, e.g. $g_{1},g_{2},\dots$ for the Gittins indices of boxes $\mathcal{B}_{1},\mathcal{B}_{2},\dots$. Finally, following along the lines of Doval (2018, Assumption 1), we make an assumption for ease of presentation.
Assumption 1.
For all boxes $\mathcal{B}=(\mathcal{D},c)$ we consider, we assume the quantities defined above satisfy $h\leq m$. One can show that this further implies [^12] These follow from the following observations and plotting $m$, $y$, and $\mathbf{E}[\max\{X,y\}]-c$ as functions of $y$. (a) $h=h^{\prime}$ is the value of $y$ such that $m=\mathbf{E}_{X\sim\mathcal{D}}[\max\{X,y\}]-c$. (b) $g$ is the value of $y$ such that $y=\mathbf{E}_{X\sim\mathcal{D}}[\max\{X,y\}]-c$. (c) As a function of $y$, the expression $\mathbf{E}_{X\sim\mathcal{D}}[\max\{X,y\}]-c$ has derivative in $[0,1]$, and at $y=0$, its value is $m-c$. $c\leq m\leq g$. If a box $\mathcal{B}$ fails this condition, we replace it with a normalized box $\mathcal{B}_{\mathrm{norm}}=(m,0)$, which trivially satisfies $h_{\mathrm{norm}}=m_{\mathrm{norm}}=g_{\mathrm{norm}}=m$.
This assumption says, roughly speaking, that every box has an open action that is sometimes worth taking. We emphasize that boxes which fail 1 do not present an obstacle. In fact, quite the opposite: the grab action is the only worthwhile action for such boxes. See Doval (2018, Section 2.2 and Appendix S.4) for further discussion. Rather than handling two cases throughout our presentation, we assume that all boxes are normalized in the way described in 1. Normalizing a box essentially removes the open action by replacing it with a “no-op”. Crucially, normalizing a box does not change the optimal surrogate value $W^{*}$, which is always $m$ for boxes that violate 1. As such, the bound we employ on the expected value of the optimal policy (Theorem E.4) remains true for the initial unnormalized system.
7.2 Failure of Local Approximation for PBOI with Rewards
Committing policies for PBOI randomize over the two actions – open and grab – and can accordingly be described by specifying the probability of the open action, call it $p$. Recalling that the surrogate value for opening is $W^{\mathrm{open}}$ and the surrogate value for grabbing is $m$, we can adapt the notion of local approximation introduced in Definition 7 for the PBOI setting as follows. [^13] Also see Scully and Doval (2024, Definition 4.2).
Definition 10 (Local $\alpha$-approximation for PBOI).
Consider a box $\mathcal{B}=(\mathcal{D},c)$. We say the a probability $p\in[0,1]$ is a local $\alpha$-approximation for $\mathcal{B}$ if for all $y$,
$(1-p)\mathbf{E}[\max\{W^{\mathrm{open}},y\}]+p\max\{m,y\}\geq\mathbf{E}[\max\{\alpha W^{*},y\}]$
Intuitively, this is saying that if we commit to grabbing the box with probability $p$ and otherwise commit to inspecting it before opening it, we get a local $\alpha$-approximation. Unfortunately, this notion of local approximation only obtains a $(1/2)$-approximation in the worst case, as proven by the following lemma.

### v2

@@ 7 The Weighing Scale Problem (Minimization)
We will now introduce the Weighing Scale (henceforth, WS) problem. A decision maker is presented with $n$ alternatives $(X_{i},c_{i})$ and a combinatorial constraint $\mathcal{F}\subseteq 2^{[n]}$; $X_{i}\geq 0$ is the random value of the alternative realized independently by a known distribution and $c_{i}\geq 0$ is a weighing cost. The only way the decision maker can determine any further information about each value $X_{i}$ beyond its distribution, is to use a weighing scale to compare it against some fixed threshold $t$ of their choosing at the additional cost of $c_{i}$ and learn whether $X_{i}\leq t$ or not. The process terminates with the decision maker selecting a feasible set of alternatives $S\subseteq\mathcal{F}$, paying their total value. Note that each alternative $(X_{i},c_{i})$ corresponds to a Costly Information MDP $\mathcal{M}_{i}$ that captures the information acquisition process described above. In particular, the available actions at the starting state of each MDP $\mathcal{M}_{i}$ are as follows:
@@
    Pick a threshold $t\in\mathrm{support}(X_{i})$ and weigh the alternative against it at a cost of $c_{i}$. Upon taking one of these actions, the MDP advances to one of two random subprocesses $\mathcal{M}_{i}^{\leq t}$ and $\mathcal{M}_{i}^{>t}$, defined over the random variables $(X_{i}|X_{i}\leq t)$ and $(X_{i}|X_{i}>t)$ respectively, based on the outcome of the weighing.
@@
From this equivalence, an instance of WS corresponds to an instance of min-CICS over the corresponding MDPs. [^11] Technically, our framework does not capture infinite horizon MDPs. However, observe that whenever the decision maker has identified that the value of the alternative lies in some interval of length $\leq c$, performing any extra weighings is suboptimal. Thus, the corresponding MDPs for the WS problem are finite horizon without loss, even for continuous random variables $X$. Our main result for this section is the following:
@@
The commitment gap of matroid-WS is at most $O(\max_{i}\kappa_{i})$, with the parameter $\kappa_{i}$ for each alternative $i\in[n]$ defined as
@@
We prove Theorem 7.1 by showing that for each alternative $i\in[n]$, the MDP $\mathcal{M}_{i}$ admits an $O(\kappa_{i})$-pointwise approximation. Furthermore, the commitment $\pi$ that achieves this condition corresponds to a very natural and efficient algorithm which we call One-Sided Halving. Finally, we note that without any assumptions on the distributions of the values, the parameters $\kappa_{i}$ can be unbounded; in particular, for heavy tail distributions, $\mu_{i}/M_{i}$ can be arbitrarily large. We show in Theorem 7.2 (proven in Appendix H) that this dependence is necessary for any pointwise approximation guarantee. Better bounds may be possible via the weaker notion of local approximation or a global argument.
@@
For any constant $\alpha\geq 1$, there exists a WS alternative that does not admit an $\alpha$-pointwise approximation.
Proof of Theorem 7.1
We begin by introducing the committing policy for WS described below, which we call the One-Sided Halving algorithm. The policy begins by weighing the alternative against some threshold $t_{2}$; if the alternative is larger, then the policy commits to no more weighings and if it is smaller, it halves its threshold from $t_{2}$ to $t_{2}/2$ and repeats, until the threshold reaches some fixed lower bound $t_{1}$.
@@
Input : An MDP $\mathcal{M}$ corresponding to a WS alternative $(X,c)$ and two thresholds $0<t_{1}<t_{2}$.
@@
Let $h$ be the solution to equation $c=\operatorname{E}\left[(X-h)^{+}\right]$. The following lemma, combined with our reduction from pointwise to local approximation (4.2) and our local approximation composition result (Theorem 4.1) directly implies Theorem 7.1.
@@
The proof of Lemma 7.3 is deferred to Appendix H. Here, we provide some intuition on how this result is obtained. A commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ for WS will declare in advance a decision tree over pre-specified thresholds against which it will weigh $X$, resulting in a Markov chain $\mathcal{M}^{\pi}$. Importantly, $\pi$ will end up partitioning the support $\mathcal{X}$ of distribution $\mathcal{D}$ into a set of intervals, corresponding to its terminal states. Furthermore, the probability of running the Markov chain $\mathcal{M}^{\pi}$ and ending up in a terminal state $t$ corresponding to some interval $I$ will be precisely $\operatorname{Pr}\left[X\in I\right]$. A pictorial representation of such Markov chains is given in Figure 7. By directly relating the water filling amortized cost on these intervals for the One-Sided Halving algorithm to the surrogate cost of the underlying MDP, we are able to obtain our pointwise approximation guarantees.
@@
8 Pandora’s Box with Optional Inspection (Maximization)
Finally, we consider the maximization version of Pandora’s Box with Optional Inspection (henceforth, PBOI), arguably the most studied Pandora’s Box variant. In PBOI, the decision-maker may select a box without opening it; we refer to this action as “grabbing” the box. There is a rich line of research addressing PBOI; we provide a detailed discussion of these results in Section 1.2. In the single-selection setting, PBOI is NP-hard but admits a PTAS (Fu et al., 2022; Beyhaghi and Cai, 2022). For matroid constraints, Beyhaghi and Kleinberg (2019) show that the commitment gap is at least $0.63$; however, their proof is nonconstructive, so the only known way to achieve this ratio is to enumerate and evaluate the exponentially many committing policies.
Currently, there is no known policy that achieves better than a $0.5$-approximation for matroid-PBOI. Additionally, this $0.5$-approximation is somewhat uninteresting. As Beyhaghi and Kleinberg (2019) observe, the optimal policy that does not grab any box will attain half of the utility from boxes that the optimal doesn’t grab, and the optimal policy that only grabs boxes will attain half of the utility from boxes that the optimal grabs. Thus, uniformly randomizing between these policies will immediately give us a $0.5$-approximation. [^12] Beyhaghi and Kleinberg (2019) make this observation in the single-item setting, but combining it with results of Singla (2017) generalizes it to the matroid setting. In this section, we present the first improvement over the trivial approximation factor of $0.5$ for matroid-max-PBOI:
@@ 8 Pandora’s Box with Optional Inspection (Maximization)
There exists an efficient randomized committing policy that achieves a $0.582$-approximation to the optimal adaptive policy for any instance of matroid-max-PBOI.
In Section 8.1 we cast PBOI into our CICS framework by considering each optional-inspection box as an MDP. A reasonable first approach would be to establish a local approximation condition for these MDPs in order to prove Theorem 8.1; however, we show that for any $\varepsilon>0$, there exist boxes for which we cannot do better than a $(0.5+\varepsilon)$-local approximation (Section 8.2). We use the intuition from these hard instances to develop a refinement of local approximation, which we call semilocal approximation; we then show that, similarly to local approximation, semilocal approximation guarantees can be composed into a (global) approximation ratio (Section 8.3). Finally, in Section 8.4 we show that this new notion of approximation suffices to establish the bound of Theorem 8.1. Omitted proofs are presented in Appendix I.
Limitations of Semilocal Approximation.
@@
There are two prices we pay in refining local approximation to semilocal approximation. First, the definition as we state it currently is specific to PBOI, though it could perhaps be generalized to other MDP families in the future. Second, our composition result for semilocal approximations only holds for matroid feasibility constraints, whereas past works showed results compatible with any frugal (roughly, “greedy”) algorithm (Singla, 2017; Gupta et al., 2019; Scully and Doval, 2024). We are not sure whether this second limitation is fundamental or might be overcome using a different proof technique.
8.1 PBOI Preliminaries
We begin by establishing the necessary notation for PBOI. All the definitions mentioned in this section were previously established for the minimization setting by Scully and Doval (2024); we simply restate them here for the maximization setting.
Fix any optional-inspection PB $\mathcal{B}=(\mathcal{D},c)$. We use $X\sim\mathcal{D}$ to denote the value of the box and $\mu=\operatorname{E}\left[X\right]$ to denote its expectation. We can model $\mathcal{B}$ as a Costly Information MDP $\mathcal{M}$ with two actions: opening the box at a cost of $c$ and transitioning to a terminal state of reward $X\sim\mathcal{D}$, and grabbing the box at a cost of 00 and transitioning to a terminal state of reward $\mu$ (Figure 8). Therefore, the set of (deterministic) local commitments for $\mathcal{M}$ consists of just two options $\{o,g\}$ and the optimality curve of $\mathcal{M}$ can be written as $f_{\mathcal{M}}(y):=\max\{f_{\mathcal{M}^{o}}(y),f_{\mathcal{M}^{g}}(y)\}$, where
$f_{\mathcal{M}^{o}}(y):=\max\{y,\operatorname{E}\left[\max(X,y)\right]-c\}\quad\text{,}\quad f_{\mathcal{M}^{g}}(y):=\max\{y,\mu\}$
are (by definition) the optimality curves under the two commitments. Using the inherent connection between optimality curves and surrogate costs, we can use these expressions to recover the following (Doval, 2018):
Definition 11 (Surrogate costs for PBOI).
Let $\mathcal{B}=(\mathcal{D},c)$ be any optional-inspection box. We define its Gittins index $g$ as the solution to equation $c=\operatorname{E}_{X\sim\mathcal{D}}\left[(X-g)^{+}\right]$ and its backup index as $h=\max(\mu,h^{\prime})$, where $h^{\prime}$ is the solution to equation $c=\operatorname{E}_{X\sim\mathcal{D}}\left[(h^{\prime}-X)^{+}\right]$. Then:
  1. 1.
    $W^{*}_{o}:=W^{*}_{\mathcal{M}^{o}}=\min\{X,g\}$ satisfies $f_{\mathcal{M}}^{o}(y)=\operatorname{E}\left[\max\{W^{*}_{o},y\}\right]$.
  2. 2.
    $W^{*}_{g}:=W^{*}_{\mathcal{M}^{g}}=\mu$ satisfies $f_{\mathcal{M}}^{g}(y)=\operatorname{E}\left[\max\{W^{*}_{g},y\}\right]$.
  3. 3.
    $W^{*}:=W^{*}_{\mathcal{M}}=\max\{W^{*}_{o},h\}$ satisfies $f_{\mathcal{M}}(y)=\operatorname{E}\left[\max\{W^{*},y\}\right]$.
These indices have an intuitive interpretation: consider the local game $(\mathcal{M},y)$ for some value of the outside option $y$. For small values of $y$, the optimal policy will immediately grab the box, and for large values of $y$ it will accept the outside option. The Gittins index $g$ corresponds to the maximum value of the outside option for which opening is preferable to stopping, and $h^{\prime}$ corresponds to the minimum value of the outside option for which opening is preferable to grabbing. In other words, the optimal policy for the local game $(\mathcal{M},y)$ will grab the box if $y\in[0,h^{\prime}]$, it will open the box if $y\in[h^{\prime},g]$ and it will accept the outside option if $y\geq g$. Note that if $g<h^{\prime}$, opening the box is never the optimal action. We refer to such boxes $\mathcal{B}=(\mathcal{D},c)$ as degenerate, and since they are never opened by any optimal policy, we can substitute them by normalized boxes $\mathcal{B}=(\mu,0)$ without loss of any generality. From now on, we will be assuming that all degenerate boxes have been normalized; this in turn implies that $h=h^{\prime}\leq\mu$ and $c\leq\mu\leq g$ (Figure 8).
Figure 8: An optional inspection box $\mathcal{B}=(\mathcal{D},c)$ can me modeled as an MDP $\mathcal{M}$ with two actions, grab and open, incuring costs 00 and $c$ and resulting to rewards $\operatorname{E}\left[X\right]$ and $X$ respectively. On the right side, we demonstrate the optimality curves for the local game $(\mathcal{M},y)$ for a box with cost $c=0.5$ and reward $X=3\cdot\mathrm{Be}(1/2)$.
8.2 Insufficiency of Local Approximation
Since any MDP $\mathcal{M}$ describing an optional-inspection box only has two actions, any (randomized) commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ is fully described by the probability $p\in[0,1]$ of committing to the grab action. Therefore, by Definition 11, the $\alpha$-local approximation condition for max-PBOI translates to
$\exists p\in[0,1]:\;\;(1-p)\cdot\operatorname{E}\left[\max\{W^{*}_{o},y\}\right]+p\cdot\operatorname{E}\left[\max\{W^{*}_{g},y\}\right]\geq\operatorname{E}\left[\max\{\alpha W^{*},y\}\right].$
We will now demonstrate a specific box where no commitment $p\in[0,1]$ can achieve a local approximation factor that is better than $1/2$.
Example 3.
For any $n\geq 1$, consider the optional-inspection box $\mathcal{B}^{n}=(\mathcal{D},c)$ where $c=n-1$ and
$\mathcal{D}=\begin{cases}1&\text{w.p. }1-\frac{1}{n^{2}}\\ n^{3}&\text{w.p. }\frac{1}{n^{2}}\\ \end{cases}.$
Straightforward computations show that the parameters of this box are $\mu=n+1-n^{-2}$, $g=n^{2}$ and $h=1+n^{2}/(n+1)$. Using these expressions and the definition of surrogate costs (Definition 11), we can proceed to write-down the local approximation condition at points $y=0$ and $y=\mu$ as follows:
Local approximation at $\text{Local approximation at $y=0$: }\;\;\alpha\leq 1-(1-p)\cdot\frac{n^{3}-n^{2}}{n^{3}+n^{2}-1}$
Local approximation at $\text{Local approximation at $y=\mu$: }\;\;\alpha\leq 1-p\cdot\frac{n^{2}-n-1+n^{-2}}{n^{2}}$
Just from these two values of $y$, we obtain that as $n\rightarrow\infty$, we have that $\alpha\leq 1-\max(p,1-p)\leq 0.5$.
To better understand this “worst-case” behavior that Example 3 demonstrates, it will be illustrative to visualize the optimality curves of the box $\mathcal{B}^{n}$ for some large value of $n$. Figure 9 shows the curves for $n=20$. We see that when $y$ is small, we have a strong preference for grabbing the box, and when $y$ is larger, we have a very slight preference for opening the box. Intuitively, we might guess that committing to the grabbing action performs quite well, since in the worst case there is only a small additive sub-optimality factor. However, local approximation requires a small multiplicative sub-optimality factor, which, as Example 3 above demonstrates, we cannot achieve.
Figure 9: The red and blue lines represent the optimality curves for grabbing $\mathcal{B}^{20}$ and opening $\mathcal{B}^{20}$, respectively. This best possible local approximation achieved for this box is $0.537$.
8.3 Semilocal Approximation
The above discussion suggests that we need to reason about additive suboptimality, not just multiplicative. We do so by introducing $(\alpha,\beta)$-semilocal approximation, defined below. Roughly speaking, the $\alpha$ captures multiplicative suboptimality, while $\beta$ captures the additive suboptimality. Formally:
Definition 12 (Semilocal Approximation).
We say that an optional-inspection box $\mathcal{B}=(\mathcal{D},c)$ admits an $(\alpha,\beta)$-semilocal approximation if there exists a probability $p\in[0,1]$ such that for all $y\in\mathbb{R}$:
$(1-p)\cdot\mathbf{E}[\max\{W^{*}_{o},y\}]+p\cdot\mathbf{E}[\max\{W^{*}_{g},y\}]\geq\mathbf{E}[\max\{\alpha W^{*},y\}]-p\beta\mu.$
Re-writing the above expression as
$(1-p)\cdot\mathbf{E}[\max\{W^{*}_{o},y\}]+p\cdot(\mathbf{E}[\max\{W^{*}_{g},y\}]+\beta\mu)\geq\mathbf{E}[\max\{\alpha W^{*},y\}],$
it becomes clear that $(\alpha,\beta)$-semilocal approximation is basically an $\alpha$-local approximation where we have boosted the value attained from grabbing by an additive factor proportional to the mean. It is important that we do this only for grabbing—boosting the value of inspecting by an additive factor would lead to poor guarantees when composing semilocal approximations in the matroid selection setting.
Equipped with the semilocal approximation definition, we now show that semilocal approximations can be composed under matroid feasibility constraints: